OpenStudy (anonymous):

2/3+3k/4=71/12 Medals!

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

Unfortunately-no.

OpenStudy (anonymous):

grr would make me more confident in my choice but find the lcm of all the denominators which would be 12 multiply all denominators that will cancel out the fractions subtract 71 by 2 then divide by 3 k=

OpenStudy (anonymous):

What does lcm stand for?

OpenStudy (anonymous):

lowest common multiple

OpenStudy (anonymous):

Would it be 12?

OpenStudy (anonymous):

yes so you would multiple all denominators by 12 numorators stays the same

OpenStudy (anonymous):

2+3x=71

OpenStudy (anonymous):

or 2+3k=71

OpenStudy (anonymous):

multiplying the denominators will cancel out the fraction

OpenStudy (anonymous):

from here subtract 2 then divide by 3

OpenStudy (anonymous):

@jdoe0001 this is correct right? don't want to be explaining wrong

OpenStudy (anonymous):

I'm sorry but you completely lost me...

OpenStudy (anonymous):

I hate math

OpenStudy (anonymous):

My result is 9k/12=756/12

jimthompson5910 (jim_thompson5910):